OpenStudy (anonymous):

Determine which ordered pairs are solutions for x - y =16. A. (8,8) B. (19,-3) C. (16,0) D. (19,3) I was thinking its C

Directrix (directrix):

If it is "c" then x = 16 and y = 0. x - y = 16 (16) - (0) = 16 16 = 16 You are correct!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Couldn't this be answer D as well as C?

Directrix (directrix):

x - y = 16 Check D. (19,3) (19) - (3) = 16 ?? 16 = 16 Yes, D is correct as is C.

Directrix (directrix):

We may need to check the other options - I was thinking the instructions were to choose ONE of the following. I was wrong.